OpenStudy (anonymous):

find the zeros for each polynomial function and give the multiplicity for each zero f(x)= -3(x+1/2)(x-4)^3

OpenStudy (whpalmer4):

Do you know how to find the zeros here?

OpenStudy (anonymous):

no do I solve for 0

OpenStudy (whpalmer4):

a polynomial can be written as a product of factors \((x-r_1)(x-r_2)...(x-r_n)\) where \(r1...r_n\) are the roots or zeros. You have \[f(x) = -3(x+\frac{1}{2})(x-4)^3\]What do you think your zeros are going to be?

OpenStudy (whpalmer4):

the zeros are the same values that make any of those product terms = 0 \[(x+\frac{1}{2}) = 0\]\[(x-4)=0\]

OpenStudy (whpalmer4):

do you know what the multiplicity of a zero is?